OpenStudy (anonymous):

Which is the distance between the point with the coordinates (-2, 3) and the line with the equation 6x-y=-3?

OpenStudy (whpalmer4):

Find the slope of line \(6x-y=-3\). Find the slope of a line perpendicular to that line (the product of the two slopes will be -1\). Now write the equation of a line with that newfound slope that goes through the point \((-2,3)\) using the point-slope formula: \[y-y_1 = m(x-x_1)\] Find the point where that line intersects with \(6x-y=-3\). You can use the distance formula to find the distance between the two points \((-2,3)\) and the point you just found: \[d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\]

OpenStudy (anonymous):

okay hold on thanks

OpenStudy (anonymous):

so it would be (-2-6)^2+(3-3)^2?

OpenStudy (whpalmer4):

No, it wouldn't be that. For starters, you didn't take the square root. Also, you aren't using the right point for the point on the line closest to \((-2,3)\). Let's check your intermediate steps. 1) what is the slope of the line \(6x-y=-3\)? 2) what is the slope of a line perpendicular to that line? 3) what is the equation of a line with the slope in 2) that goes through point \((-2,3)\)? 4) where does that line intersect with the line \(6x-y=-3\)?